Suppose that a human life is worth $6,573 and that The Woodlands has a population of $2,400. Increasing the water quality standards would reduce the risk of dying of each individual in The Woodlands from 1/63 to 1/132. What are the expected monetary benefits of the higher water quality standards? (input only numbers in your answer)
Answer #1

The expected monetary loss from a death with low quality water = 6573*1/63 = $104.33

Total population = 2400

Total expected monetary loss = 2400 * 104.33 = $250,400

The expected monetary loss from a death with better quality water = 6573*1/132 = $49.80

Total expected monetary loss = 2400*49.80 = $119,520

So, the expected monetary benefits of higher water quality = 250400 - 119520 = $130,880.

IF YOU FIND THIS ANSWER HELPFUL, PLEASE UP VOTE.